1.      An individual seller’s monthly supply of downloadable e-books is given by the equation

Qseb = −6.45 3.7Peb – 7.5W

where Qseb is number of e-books supplied each month, Peb is price of e-books in euros, and W is the hourly wage rate in euros paid by e-book sellers to workers. Assume that the price of e-books is €10.68 and the hourly wage is €10.

a.      Determine the number of e-books supplied each month.

b.     Determine the inverse supply function for an individual seller.

c.      Determine the slope of the supply curve for e-books.

Determine the new vertical intercept of the individual e-book supply curve if the hourly wage were to rise to €15 from €10.



1
Expert's answer
2021-11-29T11:32:59-0500

a.a.      Determine the market aggregate supply function.

"Qseb=\u221264.5+37.5(10.68)\u22127.5(10)=261"



b..     Determine the inverse market supply function.

Holding all other things constant, the wage rate is constant at €10, so we have

"Qseb=\u221264.5+37.5Peb\u22127.5(10)=\u2212139.5+37.5Peb"

We now solve this for Peb:

"Peb = 3.72 + 0.0267Qeb"


c. Determine the slope of the aggregate market supply curve.


when Qeb rises by one unit, Peb rises by 0.0267 euros, so the slope of the supply curve is 0.0267, which is the coefficient on Qeb in the inverse supply function. Note that it is not 37.5.